28x-2(10x-0.365)=7x+0.035

Answer by 13114(5)

28x-2(10x-0.365)=7x+0.035
28x - 20x + 0.730 = 7x + 0.035
8x + 0.730 = 7x + 0.035
8x - 7x + 0.730 = 7x - 7x + 0.035
x + 0.730 = 0.035
x + 0.730 - 0.730 = 0.035 - 0.730
x = 0.695